OpenStudy (anonymous):

Solve the Proportion. 9/24 = 12/n. @dc104

OpenStudy (anonymous):

(9/24)n = 12 n = 12/(9/24) n = 32

OpenStudy (anonymous):

Do you mind if I keep calling for you? xP @dc104

OpenStudy (anonymous):

You should be noticing a similar pattern emerging here. Move the variable (x, n, y, whatever) so that it is not in a denominator by multiplying by the variable on both sides. Divide anything that is multiplying the variable on both sides. Simplify the fraction remaining.

OpenStudy (anonymous):

In this case: 1) multiply by n on both sides 2) divide by (9/24) 3) simplify the fraction (you can do this on a calculator if needs be)

OpenStudy (anonymous):

Is it alright if I keep asking for your help? I don't want you to get frustrated with my constant pleas for help.

OpenStudy (anonymous):

I just think it's better you learn the method, as it seems to be the same each time, so you can do these yourself in future!